OpenStudy (anonymous):

Solve the equation? 3^x^3 = 9^x. I converted it to 3^(x^3) = 3^2^x Then I got x^3 = 2^x The calculator said error when I tried to find the intersection of the two graphs.

OpenStudy (anonymous):

x^3=2x not 2^x

OpenStudy (anonymous):

Ohhh ok

OpenStudy (anonymous):

really? it depends on how it is written

OpenStudy (anonymous):

\[3^{x^3}=9^x\]?

OpenStudy (anonymous):

yep that's it

OpenStudy (anonymous):

oh never mind it makes no difference. sorry. shadow is right

OpenStudy (anonymous):

\[x^3=2x\] \[x^3-2x=0\] \[x(x^2-2)=0\] \[x=0,\sqrt{2},-\sqrt{2}\]

OpenStudy (anonymous):

Yay the calc matches that! ty